### v2

@@ Abstract
We investigate the asymptotic symmetry structure of two–dimensional dilaton gravity in its $\mathcal{N}=1$ supersymmetric extension based on the $\mathfrak{osp}(1|2)$ Lie superalgebra. Within the BF theoretical framework, we analyze affine and superconformal boundary conditions for each case and systematically derive the associated asymptotic symmetry algebras. While the classical theory recovers the Virasoro algebra or its affine enhancement, the supersymmetric extension yields a classical $\mathcal{N}=1$ superconformal algebra, subject to dynamical symmetry breaking mechanisms induced by the dilaton supermultiplet. We find that the boundary behavior of the dilaton not only leads to a reduction of the full affine $\mathfrak{osp}(1|2)_{k}$ symmetry down to $\tt{O}\tt{S}p(1|2)$, but also introduces an abelian extension through commuting modes. These results reveal a novel interplay between symmetry breaking and symmetry extension in low-dimensional supergravity. Our construction generalizes previous analyses of $sl(2,\mathbb{R})$ dilaton gravity to the supersymmetric domain and offers a consistent foundation for investigating boundary dynamics beyond the Schwarzian regime.
@@ 1 Introduction
The intricate internal structure of JT gravity is fundamentally grounded in its boundary symmetry content and the formal mathematical scaffolding that supports it. The theory maintains a profound connection to conformal symmetry through its interplay with the Virasoro algebra and Schwarzian derivatives Iliesiu ; Mertens ; OvisenkoSchwarzian , and it has been systematically broadened by the implementation of relaxed boundary conditions and higher–spin extensions Gross:2016kjj ; Witten:2016iux ; Yoon . The boundary behavior can also be effectively captured using coadjoint orbit methods, which serve as powerful tools for probing symmetry–breaking phenomena and their physical ramifications Davison:2016ngz ; NarayanYoon ; Fu:2016vas ; Grumiller:2017qao ; Alkalaev:2013fsa ; Grumiller:2013swa ; Grumiller ; Kine ; Fukuyama . While the Schwarzian formalism remains a crucial link to the SYK dual, the scope of asymptotic symmetry analysis extends beyond it; in particular, the BF formulation of JT gravity offers a structurally independent yet equally robust avenue for such inquiries Grumiller:2002nm ; Banks:1990mk . Within this formalism, the dilaton field functions as a Lagrange multiplier that enforces curvature constraints, and the associated asymptotic symmetry algebra can be extended, for instance, to a Virasoro–Kac–Moody structure, via the inclusion of internal symmetry sectors Ikeda:1993fh ; Jackiw:1982hg . This flexibility underscores the utility of JT gravity as a versatile platform for studying gravitational dynamics in lower dimensions, irrespective of adherence to the Schwarzian perspective. This methodological structure aligns with the strategy developed in our previous work on bosonic $\mathfrak{sl}(3,\mathbb{R})$ JT gravity Ozer:2025bpb , and here we implement its supersymmetric analogue using the $\mathfrak{osp}(1|2)$ superalgebra. While prior analyses of JT supergravity, particularly those centered around the Schwarzian action Almheiri:2014cka ; Maldacena:2016upp ; Cadoni:1999ja ; Engelsoy:2016xyb ; Cvetic:2016eiv ; Sachdev:2019bjn ; Sachdev:2010um , focus on the boundary effective theory and its relation to the SYK model, our approach emphasizes the bulk gauge-theoretic formulation and its implications for asymptotic symmetry breaking. In contrast to the boundary-centered derivations of the ${\cal N}=1$ super–Schwarzian theory, we construct the full $\mathfrak{osp}(1|2)$ asymptotic algebra via BF theory and coadjoint orbit methods, thereby uncovering dynamical constraints absent in Schwarzian-only treatments. This strategy is conceptually aligned with our earlier $\mathfrak{sl}(3,\mathbb{R})$ construction Ozer:2025bpb , and extends the findings of Grumiller:2002nm ; Banks:1990mk ; Ikeda:1993fh ; Jackiw:1982hg into the supersymmetric domain. As such, the present work provides a structurally distinct extension of JT supergravity beyond existing formulations.
@@
In the framework of two–dimensional JT gravity, the Schwarzian formalism is frequently utilized to characterize boundary dynamics and to establish its correspondence with the SYK model Grumiller:2002nm ; Ikeda:1993fh ; Jackiw:1982hg . Nevertheless, when the objective shifts toward understanding the structure of asymptotic symmetries rather than boundary–specific features, reliance on the Schwarzian approach becomes nonessential Grumiller:2002nm ; Jackiw:1982hg . In analogy with the Chern–Simons treatment of three–dimensional gravity and the associated Brown–Henneaux boundary conditions, JT gravity admits a BF–theoretic formulation where the dilaton functions as a Lagrange multiplier enforcing curvature constraints Grumiller:2002nm ; Banks:1990mk . This perspective enables the realization of extended asymptotic symmetry algebras, such as the Virasoro or Virasoro–Kac–Moody structures, via the inclusion of internal symmetries and the imposition of compatible boundary data Ikeda:1993fh ; Jackiw:1982hg . In particular, coadjoint orbit techniques Ikeda:1993fh ; Jackiw:1982hg provide a systematic route for exploring such algebras, especially in scenarios featuring boundary–induced symmetry breaking. As a result, the BF formulation accentuates the adaptability of JT gravity in addressing holographic and asymptotic structures beyond the confines of the Schwarzian perspective.
@@
[…]
@@
Within AdS2 gravity, boundary conditions are generally classified into two dominant types: the broader affine class and the more constrained conformal class. Each class is associated with a distinct asymptotic symmetry algebra and yields unique boundary dynamics. The affine category naturally arises in the Poisson sigma model, where both the dilaton and the boundary metric are allowed to fluctuate at leading order. This framework gives rise to an infinite–dimensional centerless affine $\mathfrak{sl}(2,\mathbb{R})$ current algebra. While the Schwarzian action can be seen as a projection embedded within this more general construction, the complete theory supports a considerably richer algebraic structure. On the other hand, conformal boundary conditions impose geometric constraints that lead to a low–energy effective Schwarzian theory, describing pseudo–Goldstone modes resulting from spontaneous breaking of reparametrization symmetry. These configurations exhibit an off–shell Virasoro symmetry that reduces to ${\tt S}{\tt L}(2,\mathbb{R})$ on-shell, and are widely regarded as the gravitational duals to the infrared limit of the SYK model. Altogether, the affine and conformal classes of boundary conditions define complementary sectors in AdS2 holography and highlight the conceptual breadth of two–dimensional dilaton gravity beyond the confines of the JT framework.
@@
Unlike the well–established $sl(2,\mathbb{R})$ formulation of JT gravity, where boundary symmetries are governed by the linear Virasoro algebra, the $sl(3,\mathbb{R})$ extension offers a fertile ground for exploring more complex, nonlinear structures such as the $\mathcal{W}_{3}$ algebra Ozer:2025bpb . This framework is motivated by the pursuit of higher-spin generalizations of dilaton gravity and by the goal of understanding their impact on boundary dynamics. In particular, the inclusion of spin–3 fields alters the asymptotic symmetry structure: it not only enhances the existing symmetry but also introduces a tunable mechanism of symmetry breaking. This formulation yields fresh perspectives on how two–dimensional gravity responds to generalized boundary conditions and contributes to the broader effort to connect higher–spin JT gravity to SYK–like holographic duals. Furthermore, given that $\mathcal{W}_{3}$ algebras naturally arise in large–$N$ matrix models and in certain deformed SYK constructions, our setup may provide a plausible dual description for such deformed CFT1 theories.
@@
Recent developments have employed JT gravity models augmented with ${\tt S}{\tt L}(3,\mathbb{R})$ symmetry to explore the influence of nonlinear structures, such as the $\mathcal{W}_{3}$ algebra, on boundary dynamics Ozer:2025bpb . These investigations have clarified how higher-spin fields deform the underlying asymptotic symmetry algebra and support symmetry-breaking mechanisms. In this work, we extend the analysis beyond the purely bosonic realm by introducing a supersymmetric generalization grounded in $\mathfrak{osp}(1|2)$ supergravity. In this framework, the incorporation of a dilaton supermultiplet leads to a novel family of asymptotic field configurations. This setup not only reconfigures the algebraic structure but also furnishes a natural platform for studying supersymmetry breaking patterns and their implications for boundary dynamics.
@@
This work centers on the extension of the asymptotic symmetry algebra arising from symmetry–breaking processes, with a particular focus on the $\mathfrak{osp}(1|2)$ generalization of the JT gravity framework. This setting enables a detailed exploration of how the algebraic structure evolves and how these modifications influence the underlying physical dynamics. Previous studies have demonstrated that while a Virasoro algebra can be reconstructed within two–dimensional dilaton gravity Grumiller:2013swa , the associated conserved charges were not entirely integrable, posing challenges for the formulation of a consistent AdS2/CFT1 correspondence. This issue was subsequently resolved by introducing dilaton–dependent charges and relaxing the assumption of Casimir invariance Grumiller:2017qao , which allowed for the full manifestation of the Virasoro symmetry and reinforced the conformal field theory interpretation. These advancements highlight the critical function of extended boundary conditions in AdS2 gravity and pave the way for deeper inquiries into its holographic structure.
@@
Our construction differs conceptually and technically from conventional JT gravity formulations based on $\mathfrak{sl}(2,\mathbb{R})$ Momeni:2024ygv ; Momeni:2020tyt ; Momeni:2020zkx , particularly those that emphasize boundary actions, dilaton constraints, or effective Schwarzian descriptions. While those models often encode symmetry breaking through boundary terms and thermodynamic consistency conditions, our approach derives asymptotic symmetries directly from the bulk via the structure of the gauge field and its residual supersymmetries. This method bypasses the need for an explicit boundary action, allowing a direct algebraic characterization of the asymptotic symmetry algebra. In that sense, our work aligns more closely with generalized bulk–based symmetry analyses, such as our earlier $\mathfrak{sl}(3,\mathbb{R})$ study Ozer:2025bpb , and represents a supersymmetric extension of that framework. We do not aim to construct a complete dual quantum mechanical theory; rather, we focus on the intrinsic gauge–theoretic origin of boundary symmetries in the supersymmetric BF setting.
@@
The structure of this paper is as follows: Section 2 introduces the quantum $\mathcal{N}=1$ superconformal algebra. It begins with an overview of the conformal structure and continues with the construction of its extended version incorporating conformal spins $(\Delta,1{-}\Delta)$. Section 3 develops the holographic dictionary for $sl(2,\mathbb{R})$ JT dilaton gravity. It starts from the BF theory formulation and elaborates on its boundary realizations, including both affine and conformal boundary conditions. Section 4 focuses on the $\mathcal{N}=1$ $\mathfrak{osp}(1|2)$ extension of higher–spin dilaton gravity. It analyzes the asymptotic symmetry algebra under affine boundary conditions and then presents the superconformal boundary structure that emerges in the Drinfeld–Sokolov gauge. Section 5 summarizes the main results and discusses their theoretical and holographic significance. Section 6 acknowledges the contributors and funding sources. Finally, Appendix A provides a summary of notation, while Appendix B reviews the boundary theories and their associated symmetry structures.
@@ 2.1 Conformal structure of the quantum ${\cal N}=1$ superconformal algebra
3 The $sl(2,\mathbb{R})$ holographic dictionary
In this section, we explore the formalism of two–dimensional dilaton gravity defined on an $AdS_{2}$ background. We begin by outlining the essential features of dilaton gravity that are instrumental for constructing the holographic dictionary. This is followed by a systematic derivation of the asymptotic symmetry algebra within the $sl(2,\mathbb{R})$ framework. We then turn our attention to the role of external source terms and the resulting holographic Ward identities, which provide key insights into the structure of higher–spin black hole configurations.
@@ 3.1 JT dilaton gravity as a BF theory in two dimensions
This section provides a concise overview of higher–spin gravity in $AdS_{2}$, formulated within the framework of dilaton gravity. We adopt a gauge-theoretic perspective, treating dilaton gravity as a non–abelian theory. Specifically, our analysis focuses on the formulation of $AdS_{2}$ gravity in terms of the underlying $sl(2,\mathbb{R})$ algebraic structure.
3.2 Connection to $sl(2,\mathbb{R})$ BF theory
@@
The gauge fields $\cal A$} and $\cal X$} both take values in the Lie algebra $sl(2,\mathbb{R})$, and the trace operation $\mathfrak{tr}$ defines the bilinear form used to contract generators. To describe the dynamics of dilaton gravity in this setting, we utilize the basis of $sl(2,\mathbb{R})$ generators ${\tt L}_{i}$ with indices $i=0,\pm 1$, which obey the following commutation relations:
@@
Here, $\lambda$ is a gauge parameter taking values in the $sl(2,\mathbb{R})$ Lie algebra. These gauge transformations define the residual symmetries of the theory that preserve the asymptotic structure of the gauge fields, and they also determine how boundary symmetries emerge from bulk gauge invariance within the BF formulation. These elements provide the foundation for identifying the asymptotic symmetry algebra, which in the case of $sl(2,\mathbb{R})$ corresponds to the Virasoro algebra. The presence of the dilaton field plays a pivotal role in determining which symmetries are preserved and which are broken at the boundary. The equations of motion are given by
@@
More importantly, these gauge transformations determine the residual symmetries that persist after imposing the radial gauge near the asymptotic boundary of $\mathrm{AdS}_{2}$. Consequently, they provide the foundational structure for identifying the asymptotic symmetry algebra. The inclusion of the dilaton field $X$ is especially significant in this context: it couples to the gauge connection at the boundary and introduces dynamical constraints that restrict the admissible gauge transformations. This interaction results in a partial breaking of the full infinite–dimensional affine symmetry algebra, such as $\mathfrak{sl}(2,\mathbb{R})_{k}$ , to its corresponding finite–dimensional subalgebra, namely $\mathfrak{sl}(2,\mathbb{R})$. In this light, equations (14) and (15) do more than define local gauge redundancy; they also encapsulate the physical mechanism through which symmetry breaking is dynamically induced by the dilaton field.
@@
This representation allows for a more general class of metrics that incorporate all $sl(2,\mathbb{R})$ charges. Importantly, as long as $\delta b=0$, the specific choice of $b$ does not affect the structure of the asymptotic symmetries. This freedom enables a more versatile formulation of the metric and requires boundary conditions that preserve such generality in the gravitational context. Additionally, within the radial gauge, the reduced connection $a(t)$ is a one-form valued in the $sl(2,\mathbb{R})$ Lie algebra and is independent of the radial coordinate, i.e., $a(t)=a_{t}(t)\,\mathrm{d}t$.
Our analysis focuses on the affine boundary conditions relevant to $sl(2,\mathbb{R})$ dilaton gravity. In doing so, we make use of the methodology developed in Grumiller:2016pqb ; Ozer:2017dwk ; Ozer:2019nkv ; Ozer:2021wtx ; Ozer:2024ovo to systematically derive the associated asymptotic symmetry algebra. Based on these techniques, the most general solution consistent with asymptotically $AdS_{2}$ geometries can be expressed in the following general metric form:
@@
Accordingly, it becomes essential to specify affine $sl(2,\mathbb{R})$ boundary conditions that preserve this generalized metric structure.

### v3

@@ Abstract
We investigate the asymptotic symmetry structure of two–dimensional dilaton gravity in its $\mathcal{N}=1$ supersymmetric extension based on the $\mathfrak{osp}(1|2)$ Lie superalgebra. Within the BF theoretical framework, we analyze affine and superconformal boundary conditions and systematically derive the corresponding asymptotic symmetry algebra(ASA). While the bosonic theory reproduces the Virasoro algebra and its affine enhancement, the supersymmetric extension yields a classical $\mathcal{N}=1$ superconformal algebra whose realization is dynamically restricted by the dilaton supermultiplet. We show that the boundary behavior of the dilaton induces a controlled dynamical reduction of the full affine $\mathfrak{osp}(1|2)_{k}$ symmetry to its $\tt{O}\tt{S}p(1|2)$ stabilizer subalgebra, while simultaneously generating an abelian ideal composed of mutually commuting modes. This establishes a coherent interplay between asymptotic symmetry breaking and symmetry extension in low–dimensional supergravity. Our construction generalizes previous analyses of $\mathfrak{sl}(2,\mathbb{R})$ dilaton gravity to the supersymmetric setting and provides a consistent bulk–based framework for investigating boundary dynamics beyond the Schwarzian regime.
@@ 1 Introduction
The intricate internal structure of JT gravity is fundamentally grounded in its boundary symmetry content and the formal mathematical scaffolding that supports it. The theory maintains a profound connection to conformal symmetry through its interplay with the Virasoro algebra and Schwarzian derivatives Iliesiu ; Mertens ; OvisenkoSchwarzian , and it has been systematically broadened by the implementation of relaxed boundary conditions and higher–spin extensions Gross:2016kjj ; Witten:2016iux ; Yoon . The boundary behavior can also be effectively captured using coadjoint orbit methods, which serve as powerful tools for probing symmetry–breaking phenomena and their physical ramifications Davison:2016ngz ; NarayanYoon ; Fu:2016vas ; Grumiller:2017qao ; Alkalaev:2013fsa ; Grumiller:2013swa ; Grumiller ; Kine ; Fukuyama . While the Schwarzian formalism remains a crucial link to the SYK dual, the scope of asymptotic symmetry analysis extends beyond it; in particular, the BF formulation of JT gravity offers a structurally independent yet equally robust avenue for such inquiries Grumiller:2002nm ; Banks:1990mk . Within this formalism, the dilaton field functions as a Lagrange multiplier that enforces curvature constraints, and the associated ASA can be extended, for instance, to a Virasoro–Kac–Moody structure, via the inclusion of internal symmetry sectors Ikeda:1993fh ; Jackiw:1982hg . This flexibility underscores the utility of JT gravity as a versatile platform for studying gravitational dynamics in lower dimensions, irrespective of adherence to the Schwarzian perspective. This methodological structure aligns with the strategy developed in our previous work on bosonic $\mathfrak{sl}(3,\mathbb{R})$ JT gravity Ozer:2025bpb , and here we implement its supersymmetric analogue using the $\mathfrak{osp}(1|2)$ superalgebra.
While prior analyses of JT supergravity, particularly those centered around the Schwarzian action Almheiri:2014cka ; Maldacena:2016upp ; Cadoni:1999ja ; Engelsoy:2016xyb ; Cvetic:2016eiv ; Sachdev:2019bjn ; Sachdev:2010um , focus on the boundary effective theory and its relation to the SYK model, our approach emphasizes the bulk gauge-theoretic formulation and its implications for asymptotic symmetry breaking. In contrast to the boundary-centered derivations of the ${\cal N}=1$ super–Schwarzian theory, we construct the full $\mathfrak{osp}(1|2)$ asymptotic algebra via BF theory and coadjoint orbit methods, thereby uncovering dynamical constraints absent in Schwarzian-only treatments. This strategy is conceptually aligned with our earlier $\mathfrak{sl}(3,\mathbb{R})$ construction Ozer:2025bpb , and extends the findings of Grumiller:2002nm ; Banks:1990mk ; Ikeda:1993fh ; Jackiw:1982hg into the supersymmetric domain. As such, the present work provides a structurally distinct extension of JT supergravity beyond existing formulations.
@@
In the framework of two–dimensional JT gravity, the Schwarzian formalism is frequently utilized to characterize boundary dynamics and to establish its correspondence with the SYK model Grumiller:2002nm ; Ikeda:1993fh ; Jackiw:1982hg . Nevertheless, when the objective shifts toward understanding the structure of asymptotic symmetries rather than boundary–specific features, reliance on the Schwarzian approach becomes nonessential Grumiller:2002nm ; Jackiw:1982hg . In analogy with the Chern–Simons treatment of three–dimensional gravity and the associated Brown–Henneaux boundary conditions, JT gravity admits a BF–theoretic formulation where the dilaton functions as a Lagrange multiplier enforcing curvature constraints Grumiller:2002nm ; Banks:1990mk . This perspective enables the realization of extended ASA’s, such as the Virasoro or Virasoro–Kac–Moody structures, via the inclusion of internal symmetries and the imposition of compatible boundary data Ikeda:1993fh ; Jackiw:1982hg . In particular, coadjoint orbit techniques Ikeda:1993fh ; Jackiw:1982hg provide a systematic route for exploring such algebras, especially in scenarios featuring boundary–induced symmetry breaking. As a result, the BF formulation accentuates the adaptability of JT gravity in addressing holographic and asymptotic structures beyond the confines of the Schwarzian perspective.
@@
This phrasing, however, is not intended to suggest a replacement or extension of the Schwarzian theory at the level of physical content or dynamical observables. Instead, our approach focuses on a methodological generalization: we continue the asymptotic symmetry analysis in the BF formalism, now within the $\mathfrak{osp}(1|2)$ supersymmetric setting. This construction is conceptually aligned with our earlier $\mathfrak{sl}(3,\mathbb{R})$ study Ozer:2025bpb , wherein the coadjoint orbit framework and bulk gauge formulation were used to realize symmetry-breaking mechanisms. The present work provides a supersymmetric parallel to that framework and should be seen as a structural, rather than physical, extension beyond the Schwarzian perspective.
@@
Within AdS2 gravity, boundary conditions are generally classified into two dominant types: the broader affine class and the more constrained conformal class. Each class is associated with a distinct ASA and yields unique boundary dynamics. The affine category naturally arises in the Poisson sigma model, where both the dilaton and the boundary metric are allowed to fluctuate at leading order. This framework gives rise to an infinite–dimensional centerless affine $\mathfrak{sl}(2,\mathbb{R})$ current algebra. While the Schwarzian action can be seen as a projection embedded within this more general construction, the complete theory supports a considerably richer algebraic structure. On the other hand, conformal boundary conditions impose geometric constraints that lead to a low–energy effective Schwarzian theory, describing pseudo–Goldstone modes resulting from spontaneous breaking of reparametrization symmetry. These configurations exhibit an off–shell Virasoro symmetry that reduces to ${\tt S}{\tt L}(2,\mathbb{R})$ on-shell, and are widely regarded as the gravitational duals to the infrared limit of the SYK model. Altogether, the affine and conformal classes of boundary conditions define complementary sectors in AdS2 holography and highlight the conceptual breadth of two–dimensional dilaton gravity beyond the confines of the JT framework.
@@
Unlike the well–established $\mathfrak{sl}(2,\mathbb{R})$ formulation of JT gravity, where boundary symmetries are governed by the linear Virasoro algebra, the $\mathfrak{sl}(3,\mathbb{R})$ extension offers a fertile ground for exploring more complex, nonlinear structures such as the $\mathcal{W}_{3}$ algebra Ozer:2025bpb . This framework is motivated by the pursuit of higher-spin generalizations of dilaton gravity and by the goal of understanding their impact on boundary dynamics. In particular, the inclusion of spin–3 fields alters the asymptotic symmetry structure: it not only enhances the existing symmetry but also introduces a tunable mechanism of symmetry breaking. This formulation yields fresh perspectives on how two–dimensional gravity responds to generalized boundary conditions and contributes to the broader effort to connect higher–spin JT gravity to SYK–like holographic duals. Furthermore, given that $\mathcal{W}_{3}$ algebras naturally arise in large–$N$ matrix models and in certain deformed SYK constructions, our setup may provide a plausible dual description for such deformed CFT1 theories.
@@
Recent developments have employed JT gravity models augmented with ${\tt S}{\tt L}(3,\mathbb{R})$ symmetry to explore the influence of nonlinear structures, such as the $\mathcal{W}_{3}$ algebra, on boundary dynamics Ozer:2025bpb . These investigations have clarified how higher-spin fields deform the underlying ASA and support symmetry-breaking mechanisms. In this work, we extend the analysis beyond the purely bosonic realm by introducing a supersymmetric generalization grounded in $\mathfrak{osp}(1|2)$ supergravity. In this framework, the incorporation of a dilaton supermultiplet leads to a novel family of asymptotic field configurations. This setup not only reconfigures the algebraic structure but also furnishes a natural platform for studying supersymmetry breaking patterns and their implications for boundary dynamics.
@@
This work centers on the extension of the ASA arising from symmetry–breaking processes, with a particular focus on the $\mathfrak{osp}(1|2)$ generalization of the JT gravity framework. This setting enables a detailed exploration of how the algebraic structure evolves and how these modifications influence the underlying physical dynamics. Previous studies have demonstrated that while a Virasoro algebra can be reconstructed within two–dimensional dilaton gravity Grumiller:2013swa , the associated conserved charges were not entirely integrable, posing challenges for the formulation of a consistent AdS2/CFT1 correspondence. This issue was subsequently resolved by introducing dilaton–dependent charges and relaxing the assumption of Casimir invariance Grumiller:2017qao , which allowed for the full manifestation of the Virasoro symmetry and reinforced the conformal field theory interpretation. These advancements highlight the critical function of extended boundary conditions in AdS2 gravity and pave the way for deeper inquiries into its holographic structure.
@@
Our construction differs conceptually and technically from conventional JT gravity formulations based on $\mathfrak{sl}(2,\mathbb{R})$ Momeni:2024ygv ; Momeni:2020tyt ; Momeni:2020zkx , particularly those that emphasize boundary actions, dilaton constraints, or effective Schwarzian descriptions. While those models often encode symmetry breaking through boundary terms and thermodynamic consistency conditions, our approach derives asymptotic symmetries directly from the bulk via the structure of the gauge field and its residual supersymmetries. This method bypasses the need for an explicit boundary action, allowing a direct algebraic characterization of the ASA. In that sense, our work aligns more closely with generalized bulk–based symmetry analyses, such as our earlier $\mathfrak{sl}(3,\mathbb{R})$ study Ozer:2025bpb , and represents a supersymmetric extension of that framework. We do not aim to construct a complete dual quantum mechanical theory; rather, we focus on the intrinsic gauge–theoretic origin of boundary symmetries in the supersymmetric BF setting.
@@
The structure of this paper is as follows: Section 2 introduces the quantum $\mathcal{N}=1$ superconformal algebra. It begins with an overview of the conformal structure and continues with the construction of its extended version incorporating conformal spins $(\Delta,1{-}\Delta)$. Section 3 develops the holographic dictionary for $\mathfrak{sl}(2,\mathbb{R})$ JT dilaton gravity. It starts from the BF theory formulation and elaborates on its boundary realizations, including both affine and conformal boundary conditions. Section 4 focuses on the $\mathcal{N}=1$ $\mathfrak{osp}(1|2)$ extension of higher–spin dilaton gravity. It analyzes the ASA under affine boundary conditions and then presents the superconformal boundary structure that emerges in the Drinfeld–Sokolov gauge. Section 5 summarizes the main results and discusses their theoretical and holographic significance. Section 6 acknowledges the contributors and funding sources. Finally, Appendix A provides a summary of notation, while Appendix B reviews the boundary theories and their associated symmetry structures.
@@ 2.1 Conformal structure of the quantum ${\cal N}=1$ superconformal algebra
3 The $\mathfrak{sl}(2,\mathbb{R})$ holographic dictionary
In this section, we explore the formalism of two–dimensional dilaton gravity defined on an $AdS_{2}$ background. We begin by outlining the essential features of dilaton gravity that are instrumental for constructing the holographic dictionary. This is followed by a systematic derivation of the ASA within the $\mathfrak{sl}(2,\mathbb{R})$ framework. We then turn our attention to the role of external source terms and the resulting holographic Ward identities, which provide key insights into the structure of higher–spin black hole configurations.
@@ 3.1 JT dilaton gravity as a BF theory in two dimensions
This section provides a concise overview of higher–spin gravity in $AdS_{2}$, formulated within the framework of dilaton gravity. We adopt a gauge-theoretic perspective, treating dilaton gravity as a non–abelian theory. Specifically, our analysis focuses on the formulation of $AdS_{2}$ gravity in terms of the underlying $\mathfrak{sl}(2,\mathbb{R})$ algebraic structure.
3.2 Connection to $\mathfrak{sl}(2,\mathbb{R})$ BF theory
@@
The gauge fields $\cal A$} and $\cal X$} both take values in the Lie algebra $\mathfrak{sl}(2,\mathbb{R})$, and the trace operation $\mathfrak{tr}$ defines the bilinear form used to contract generators. To describe the dynamics of dilaton gravity in this setting, we utilize the basis of $\mathfrak{sl}(2,\mathbb{R})$ generators ${\tt L}_{i}$ with indices $i=0,\pm 1$, which obey the following commutation relations:
@@
Here, $\lambda$ is a gauge parameter taking values in the $\mathfrak{sl}(2,\mathbb{R})$ Lie algebra. These gauge transformations define the residual symmetries of the theory that preserve the asymptotic structure of the gauge fields, and they also determine how boundary symmetries emerge from bulk gauge invariance within the BF formulation. These elements provide the foundation for identifying the ASA, which in the case of $\mathfrak{sl}(2,\mathbb{R})$ corresponds to the Virasoro algebra. The presence of the dilaton field plays a pivotal role in determining which symmetries are preserved and which are broken at the boundary. The equations of motion are given by
@@
More importantly, these gauge transformations determine the residual symmetries that persist after imposing the radial gauge near the asymptotic boundary of $\mathrm{AdS}_{2}$. Consequently, they provide the foundational structure for identifying the ASA. The inclusion of the dilaton field $X$ is especially significant in this context: it couples to the gauge connection at the boundary and introduces dynamical constraints that restrict the admissible gauge transformations. This interaction results in a partial breaking of the full infinite–dimensional affine symmetry algebra, such as $\mathfrak{sl}(2,\mathbb{R})_{k}$ , to its corresponding finite–dimensional subalgebra, namely $\mathfrak{sl}(2,\mathbb{R})$. In this light, equations (14) and (15) do more than define local gauge redundancy; they also encapsulate the physical mechanism through which symmetry breaking is dynamically induced by the dilaton field.
@@
This representation allows for a more general class of metrics that incorporate all $\mathfrak{sl}(2,\mathbb{R})$ charges. Importantly, as long as $\delta b=0$, the specific choice of $b$ does not affect the structure of the asymptotic symmetries. This freedom enables a more versatile formulation of the metric and requires boundary conditions that preserve such generality in the gravitational context. Additionally, within the radial gauge, the reduced connection $a(t)$ is a one-form valued in the $\mathfrak{sl}(2,\mathbb{R})$ Lie algebra and is independent of the radial coordinate, i.e., $a(t)=a_{t}(t)\,\mathrm{d}t$.
Our analysis focuses on the affine boundary conditions relevant to $\mathfrak{sl}(2,\mathbb{R})$ dilaton gravity. In doing so, we make use of the methodology developed in Grumiller:2016pqb ; Ozer:2017dwk ; Ozer:2019nkv ; Ozer:2021wtx ; Ozer:2024ovo to systematically derive the associated ASA. Based on these techniques, the most general solution consistent with asymptotically $AdS_{2}$ geometries can be expressed in the following general metric form:
@@
Accordingly, it becomes essential to specify affine $\mathfrak{sl}(2,\mathbb{R})$ boundary conditions that preserve this generalized metric structure.